OpenStudy (anonymous):

Find the inclination θ (in radians AND degrees) of a line passing through the points: (2,-2) (7,9)

OpenStudy (anonymous):

slope=tan theta

OpenStudy (anonymous):

slope=9--2/7-2=11/5

OpenStudy (anonymous):

theta=tan inverse(11/5)

OpenStudy (anonymous):

how would you convert 11/5 into degrees and radians?

OpenStudy (anonymous):

65.55 degrees

OpenStudy (anonymous):

yeah how do you convert it ?

OpenStudy (anonymous):

finding the slope is easy. What I am not aware of as of yet is how the conversion process works.

OpenStudy (anonymous):

i will tell how

OpenStudy (anonymous):

@Krishnadas can you explain?

OpenStudy (anonymous):

X radians=X degree*pi/180

OpenStudy (anonymous):

X degrees=X radians*180/pi

OpenStudy (anonymous):

like 30 degrees=30*pi/180=pi/6

OpenStudy (anonymous):

got that??